# Gaussian Control with Hierarchical Semantic Graphs in 3D Human Recovery

### v1

@@ Abstract
Although 3D Gaussian Splatting (3DGS) has recently made progress in 3D human reconstruction, it primarily relies on 2D pixel-level supervision, overlooking the geometric complexity and topological relationships of different body parts. To address this gap, we introduce the Hierarchical Graph Human Gaussian Control (HUGS) framework for achieving high-fidelity 3D human reconstruction. Our approach involves leveraging explicitly semantic priors of body parts to ensure the consistency of geometric topology, thereby enabling the capture of the complex geometrical and topological associations among body parts. Additionally, we disentangle high-frequency features from global human features to refine surface details in body parts. Extensive experiments demonstrate that our method exhibits superior performance in human body reconstruction, particularly in enhancing surface details and accurately reconstructing body part junctions. Codes are available at https://wanghongsheng01.github.io/HUGS/.
@@
UTF8gbsn 22footnotetext: Corresponding Author.
@@ 1 Introduction
With the development of technologies such as computer graphics, virtual reality, and artificial intelligence, the generation of articulate 3D digital humans has become a popular research field. It holds immense potential value in industries like holographic transmission, and metaverse. Traditional 3D representation methods [24, 5, 17, 13, 50, 41, 9] like meshes and point clouds require dense views for human reconstruction [6, 15, 46, 19, 3, 47, 33, 16], limiting their applications in sparse views for human reconstruction.
Recently, 3D Gaussian splatting (3DGS) [20], an efficient point-based 3D representation method, has been proposed for high-fidelity rendering quality and real-time rendering speed. Recent 3DGS works on human reconstruction generally rely on 2D pixel-level supervision to optimize initial coarse Gaussian point distributions [20]. However, this approach ignores the geometrical complexity of body parts and the motion correlations between different body parts. During human motion, local occlusion may occur, resulting in semantic ambiguity at the junction of body parts, thereby causing local geometric distortions. Simultaneously, complex motion deformations bring about significant local structural discrepancies on the body surface, which usually reflect important details such as clothing wrinkles. Pixel-level supervision typically smoothens the 3D discrepancies during complex motion deformations, leading to the loss and blurring of important details. As such, learning the complex geometry and kinematic relationships of body parts is demanding in 3DGS human reconstruction.
To bridge the gap, we propose an effective framework Human Gaussian Control with Hierarchical Semantic Graphs (HUGS). HUGS incorporates the Semantic kinematic topology module, learning semantic attributes consistency and kinematic topological association of each Gaussian to capture the complex geometric structures and the linkage effects between different body parts. Furthermore, surface details of significant discrepancies primarily occur in high-frequency regions in body parts. Therefore, based on semantic priors and topological association of the Gaussian points, Surface Disentanglement module disentangle high-frequency features from human global features and densify corresponding Gaussian points, to refine surface details of significant discrepancies such as clothing wrinkles.
@@
    We propose a Gaussian human reconstruction method control with hierarchical semantic graphs, effectively addressing the problem of local occlusions, which cause semantic information confusion at the junctions of body parts, thus more accurately reconstructing the geometric structure and surface details of the human body.
@@
    We introduce a surface disentanglement method that combines semantic and topological information, disentangling high-frequency features from global human features. To guide the refinement of the density and distribution of Gaussian points in local structures with significant disparities and better restore the surface detail features.
@@
    To validate the effectiveness of our method, we conduct extensive experiments and comparisons. Compared to existing methods, the results demonstrate significant improvements in rendering quality and body structure, especially in the surface details and junction of body parts.
@@ 2 Related Work
2.1 NeRF on Human Body
NeRF (Neural Radiance Field [29]) is an implicit representation method known for high-quality 3D reconstruction and rendering. Due to its generalizable features and ray-sampling rendering capabilities, NeRF has found applications in various domains [38, 4, 49, 28, 2, 1, 45, 48, 7, 22, 40]. Recently, researchers have begun exploring NeRF’s potential for human body reconstructio [42, 44, 31, 34]. However, existing works applying NeRF to the human body exhibit certain limitations. For instance, skeleton-based NeRF generative models struggle to capture non-rigid deformations [31]. Many of these models are designed for specific scenes, necessitating significantly longer training times [42], prior information about performers [31], or encountering challenges with multi-person scenes leading to reconstruction failures [44]. The inherent implicit representation of NeRF limits its ability to accurately represent fine human appearance details and subtle pose changes [44, 31]. Moreover, the computational demands of complex ray-tracing and neural networks make it challenging to apply NeRF industrially. In contrast, the recently introduced 3D Gaussian splatting method [20] addresses NeRF’s shortcomings in 3D human reconstruction and rendering.
@@ 2.1 NeRF on Human Body
3D Gaussian Splatting [20] is an explicit scene representation technique utilizing flexible Gaussian ellipsoids, enabling more versatile 3D scene reconstruction and faster rendering compared to NeRF. Human Gaussian Splatting [30] has highlighted 3D Gaussian Splatting as an efficient alternative to NeRF. This study focuses on the application of 3D Gaussian Splatting in human body reconstruction [14, 25, 23, 26, 21, 30]. While extracting the human body from multi-view videos using 3D Gaussian Splatting is feasible [30, 26], challenges persist in reconstructing high-frequency regions with high quality [30]. The randomness in Gaussian generation during reconstruction can hinder the extraction of 3D human body mesh information [14] and lead to excessive reliance on precise pose inputs [26]. Insufficient human body geometry information may result in parts not properly following human movements (e.g., cloth deformation with poses) [30]. We propose that these issues can be mitigated through supervised Gaussian generation with human semantics, enhancing the recovery of detailed information such as cloth wrinkles from monocular view videos [14].
@@ 3 Methods
We propose Human Gaussian Control with Hierarchical Semantic Graphs (HU-GS) to generate Gaussian humans, ensuring both realistic human appearance and anatomical structure. To learn the semantic and topological association of body parts, we introduce Semantic Kinematic Topology in Section 3.1, aligning human semantic and kinematic topology with 3D humans to capture complex geometric features of body parts and kinematic correlations. In Section 3.2, we introduce Surface Disentanglement, which disentangles high-frequency features from global human body features, refining the local structures of significant discrepancies on the human surface.
Figure 2: We introduce Human Gaussian Control with Hierarchical Semantic Graphs (HUGS) as a method for generating Gaussian humans, ensuring both realistic human appearance and anatomical structure. To better understand the semantic and topological association of body parts, we present the concept of Semantic Kinematic Topology in Section 3.1, aligning human semantic and kinematic topology with 3D humans to capture complex geometric features of body parts and kinematic correlations. In Section 3.2, we propose Surface Disentanglement, aiming to disentangle high-frequency features from global human body features, thereby refining the local structures of significant discrepancies on the human surface.
@@ 3.1 Preliminary
SMPL [27]: SMPL is a pre-trained parametric human model representing body shape and pose. In SMPL, body shape and pose are controlled by two main parameters: $\beta$ and $\theta$, where $\beta$ represents 10 parameters for proportions such as height, weight, and head-to-body ratio, and $\theta$ represents 75 parameters for overall body motion posture and relative angles of 24 joints. These parameters enable the SMPL model to generate various human body shapes and poses.
@@
3.2 Semantic Kinematic Topology
Applying 3D Gaussian Splatting directly to human reconstruction had several challenges. Due to relying on 2D pixel-level supervision to refine the initial coarse Gaussian points distribution. This approach lacks explicit learning of Gaussian points according to various body parts of the human anatomy and ignores the topological relationships between these parts. The human complex motion may bring about local structures being occluded, generating semantic confusion in Gaussian points at the junction of body parts. Gaussian points fail to correctly match their corresponding spatial positions in the human body, resulting in geometric distortion in the occluded parts. Therefore, considering the intrinsic geometric structures of different body parts and their topological relationships is crucial in addressing the reconstruction problem of local geometric distortion in the occluded parts.
@@ 3.2 Semantic Kinematic Topology
Semantic Alignment Gaussian: The semantic attributes of Gaussian points explicitly inject semantic information into the optimization process of 3D Gaussians. Leveraging semantic information allows us to independently optimize specific geometric structures for each body part. Therefore, we explore rendering attributes correlated with corresponding body parts by aligning the geometric structure of each body part with its corresponding semantics. To achieve this, we propose a semantic consistency constraint to guide the independent optimization of each body part. The loss function formula corresponding to semantic consistency constraint is shown as follows.
$\mathcal{L}_{\text{semantic}}=\frac{1}{N}\sum_{i=1}^{N}(D(\mathcal{X}_{i},y_{i})+\frac{1}{|Q_{i}|}\sum_{j\in Q_{i}}D(\mathcal{X}_{j},y_{i})).$ (3)
where $N$ represents the total number of Gaussian points, $y_{i}$ represents the semantic attribute of Gaussian point $i$, $Q_{i}$ contains the $k$ nearest neighbors of Gaussian point $i$ in 3D space, not 2D-pixel space. $\mathcal{X}_{i}$ represents the semantic information of the pixel rendered by Gaussian point $i$, 3DGS employs point-based rendering techniques ($\alpha$-blending) to render the color attributes of 3D Gaussian points onto the 2D-pixel plane, function $D$($\cdot$) is used to determine whether the Gaussian point semantic properties are consistent with the true semantics corresponding to the pixels to which they are mapped. Similarly, we blend the semantic attributes of the human body from 3D space to a 2D-pixel plane by mixing them in the order of the influence of on pixels.The rendering formula is as follows:
@@
where $\mathcal{X}_{i}$ represents the semantic labels of pixel $i$, which are derived from Gaussian point semantic attribute $\alpha$-blending, $\mathcal{N}$ represents the set of all Gaussian points that overlap with the pixels rendered by Gaussian point $i$, $y_{p}$ represents the semantic attribute corresponding to the 3D Gaussian point $p$, and $\alpha_{p}$ represents the influence factor of the Gaussian point $p$ for rendering pixels.
Inter-Joint Graph: While the semantic consistency constraint independently optimizes the geometric structure of each body part, different body parts of the human body are not entirely independent. Various body parts have mutual influences and collaborative effects on each other. For example, the position and posture of the arms may be influenced by the positions and postures of the palms and shoulders. However, due to the complexity of human motion, the linkage effects between body parts cannot be explicitly expressed. Therefore, we explore using the explicit topological structure contained in 3D Gaussians to learn the implicit linkage effects among body parts. To achieve this, we introduce topological coherence constraints to learn the kinematic correlations between different body parts.
To incorporate topological information into the human reconstruction process, we utilize the spatial position information of Gaussian points to create a topological graph for the human, where each Gaussian point corresponds to a node in the graph. The attributes of the nodes can include information such as the color and transparency of the Gaussian points. By computing the distances between Gaussian points and setting a distance threshold, we establish edge relationships between nodes, with weights represented by the distances. The process of constructing edges can be expressed as follows:
$E=\left\{\left(p_{i},p_{j}\right)\mid\left\|x_{i}-x_{j}\right\|_{2}<r\right\}.$ (5)
where $E$ represents the adjacency matrix of the human topological graph, $x_{i}$ and $x_{j}$ represent the 3D position information of Gaussian points corresponding to nodes $p_{i}$ and $p_{j}$ respectively, $r$ represents the maximum threshold of spatial distance between Gaussian points $p_{i}$ and $p_{j}$ to determine whether there exists an edge between nodes $x_{i}$ and $x_{j}$ in the topological graph.
@@
In the topological graph, information propagation and sampling are conducted through a random walk algorithm [10, 37, 39, 32] to capture the topological association between different body parts. We construct contrast learning samples that combine topological associations with prior knowledge of collaborative effects among body parts in the human kinematic tree, in order to learn topological coherence constraints among body parts. The loss function formula corresponding to the topological coherence constraint is shown as follows:
where $N$ represents the total number of Gaussian points, $M$ represents the number of positive and negative samples used for contrast learning, $t_{i}$ represents the embedding vector for node $i$, which contains topological information obtained by the random walk algorithm, $t_{j}^{+}$ represent the positive samples from body parts that have a linkage effect with the body part corresponding to node $i$, such as Arm and Hand, $t_{j}^{-}$ is the negative sample from body parts that do not have a linkage effect with the body part corresponding to node $i$.
3.3 Surface Disentanglement
Based on Semantic Kinematic Topology, we initially express both the geometric structure of the body parts and the topological correlations between them. However, we observed fuzzy geometric shapes in the local structures with significant discrepancies on the human surface, such as clothing wrinkles and muscle textures. These discrepancies are primarily found in high-frequency regions of the human surface. Therefore, capturing high-frequency features becomes essential to guide the optimization of local structures with significant differences. As a result, we explored disentangling high-frequency features from the global human features and utilizing them to refine Gaussian point density and distribution corresponding to the local structures with significant discrepancies.
To address this challenge, we introduce the Surface Disentanglement Module. High-frequency signals manifest in the human topological graph as outlier nodes that significantly differ from neighboring nodes. We cluster the human topological graph into $k$ distinct clusters based on node attributes, topological features, and semantic attributes. In each iteration, we identify specific nodes within a cluster as candidate high-frequency nodes by evaluating the average magnitude of structural discrepancies between the selected node and all other nodes within the same cluster. Ultimately, we select the node with the highest average magnitude of structural discrepancies within each cluster as the high-frequency node. The algorithmic formula is as follows:
$C_{1},C_{2},\ldots,C_{k}=\text{Cluster}(A_{i},T_{i},S_{i}).$ (7)
$H_{k}=\mathop{\mathrm{argmax}}\limits_{i\in C_{k}}\frac{1}{|C_{k}|-1}\sum_{j\in C_{k},j\neq i}\text{similar}\left(A_{i},A_{j}\right).$ (8)
where $A_{i}$ represents the basic attribute of Gaussian points (color, opacity, and radii), $T_{i}$ represents the topological feature of Gaussian points obtained through random walk algorithm, $S_{i}$ represents the semantic attribute of Gaussian points, $\text{Cluster}(\cdot)$ is the K-means node clustering algorithm, $C_{k}$ represents the $k$ cluster after clustering, $\text{similar}(\cdot)$ is used to evaluate the similarity between two nodes.

### v3

@@ Abstract
Although 3D Gaussian Splatting (3DGS) has recently made progress in 3D human reconstruction, it overlooked blurred details in the reconstruction of the human body, particularly at the junctions and surface features. To address this gap, we introduce the Hierarchical Graph Human Gaussian Control (HUGS) framework for achieving high-fidelity 3D human reconstruction. Our method enhances the learning of relationships in motion at junctions of the human body by innovatively establishing semantic-level constraints. Additionally, we improve the learning of internal appearance connections within different parts of the human body by extracting high-frequency information between Gaussian points with identical semantic labels, thereby reconstructing high-frequency details within each part. Extensive experiments demonstrate that our method excels in human body reconstruction, especially at junctions and surface features.Codes are available at https://wanghongsheng01.github.io/HUGS/.
@@
UTF8gbsn 22footnotetext: Corresponding Author. This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
With the development of technologies such as computer graphics, virtual reality, and artificial intelligence, the generation of articulate 3D digital humans has become a popular research field. It holds immense potential value in industries like holographic transmission, and the metaverse. Traditional 3D representation methods [24, 3, 14, 10, 47, 39, 6] like meshes and point clouds require dense views for human reconstruction [4, 12, 44, 18, 2, 45, 33, 13], limiting their applications in sparse views for human reconstruction.
Current research incorporates the SMPL model into the 3DGS process [19, 46, 16, 23]. However, these methods still suffer from blurred details in the reconstruction of the human body, particularly at the junctions and surface features. These issues arise because existing approaches only inherit positional information from the SMPL model while neglecting the inherent connections between different body parts. Additionally, in monocular dynamic scenarios, the existing 2D pixel-level supervision often smooths out 3D discrepancies during complex motion deformations, failing to deeply capture the intricate relationships between local Gaussian points within a region. This leads to the blurring of important details such as clothing wrinkles and muscle textures.
To bridge the gap, we propose Human Gaussian Control with Hierarchical Semantic Graphs (HUGS), to deeply learn the relationships between Gaussian points both at the connections of different body parts and within body parts. To address the blurring issues at body part junctions in 3D reconstruction, we introduce the Inter-Semantic Kinematic Topology module, which improves the learning of relationships in motion at these junctions. In this module, we innovatively inherit semantic information from SMPL into Gaussian points to establish semantic-level constraints. Additionally, we construct a 3D Gaussian point graph structure and employ a random walk method to provide positional embedding for Gaussian points, thereby creating contrastive learning samples based on the prior human topological Hierarchical structure. Meanwhile, we propose the Intro-Surface Disentanglement module to handle the blurriness in every internal body part. By extracting high-frequency information between Gaussian points with the same semantic label based on a graph structure, we subsequently increase the density of Gaussian points at these high-frequency locations.
@@
    We propose an Inter-Semantic Kinematic Topology method based on a topological graph of Gaussian point semantic relationships. This method addresses the blurring issues at the connections of different body parts in 3D reconstruction by learning the relationships in motion at the junctions.
@@
    We introduce an Intro-Surface Disentanglement method based on high-frequency extraction between Gaussian points with the same semantics, structured as a graph. This approach handles the internal blurriness within body parts by learning the relationships in appearance.
@@
    To validate the effectiveness of our method, we conduct extensive experiments and comparisons. Compared to existing methods, the results demonstrate significant improvements in rendering quality and body structure, especially at the junctions and surface features.
@@ 2 Related Work
2.1 Statistical Parametric Human Body Model
The Statistical Parametric Human Body Model (SMPL) is a parametric model capable of accurately representing human body shape and pose[27]. One approach utilized SMPL for body pose and shape recovery, leveraging the optimization results generated by SMPLify as supervision signals to train an end-to-end network, thereby enhancing the realism of the reconstruction results[1]. Building on this, another approach trained 91 keypoint detectors. Subsequently, they optimized the SMPL model parameters to fit these keypoints and further proposed a random forest regression method to directly regress these parameters, albeit at the expense of accuracy[21]. Some researchers employed synthesized data rendered to train a fully convolutional model for depth and body part segmentation[41]. In another work, 2D and 3D joint positions were estimated by fitting a bound skeleton model. While 3D rotations could be recovered post-optimization, this approach directly outputs rotations and surface vertices from images[30]. Another method regressed SMPL pose, shape, and camera parameters, using synthetic data for supervised training and unsupervised fine-tuning during inference, with 2D joints, contours, and motion losses.[40].One study regressed SMPL and camera parameters and applied adversarial training, where the predicted results were input to a discriminator network for real-vs-fake classification[17]. A multi-task approach was proposed to estimate 2D/3D pose, pixel segmentation, and volumetric shape, utilizing real volumetric data generated by SMPL, but without embedding the SMPL function within the network.[42]. A deep learning method named SMPLR was subsequently introduced. It accurately recovers 3D human pose and shape from RGB images by estimating 3D joints and sparse landmarks, followed by the parameter regression[29].Previous studies utilizing SMPL for human body reconstruction overlooked the relationships in motion at the junctions, leading to blurring issues at these connection points. To address this, we propose transferring the semantic information from SMPL to Gaussian points. This approach will establish semantic-level constraints, effectively resolving the blurring issues at the junctions during reconstruction.
@@ 2.1 Statistical Parametric Human Body Model
3D Gaussian Splatting[19] is an explicit scene representation that technique utilizing flexible Gaussian ellipsoids, enabling more versatile 3D scene reconstruction and faster rendering compared to NeRF. Human Gaussian Splatting[31] has showcased 3D Gaussian Splatting as an efficient alternative to NeRF. This study focuses on the application of 3D Gaussian Splatting in human body reconstruction[11, 25, 22, 26, 20, 31]. While extracting the human body from multi-view videos using 3D Gaussian Splatting is feasible[31, 26], challenges persist in reconstructing high-frequency regions within various parts of the human body with high quality[31]. The randomness in Gaussian generation during reconstruction can hinder the extraction of 3D human body mesh information[11] and lead to an excessive reliance on precise pose inputs[26]. Insufficient human body geometry information may result in parts not properly following human movements (e.g., cloth deformation with poses)[31]. We propose that these blurry details within parts of the human body can be mitigated through supervised Gaussian generation that incorporates human semantics. By extracting high-frequency information between Gaussian points with the same semantic label, we can learn the internal relationships in appearance within various parts of the human body[11].
@@ 3 Methods
We propose Human Gaussian Control with Hierarchical Semantic Graphs (HUGS) to generate Gaussian humans, ensuring both junctions of human body and realistic human appearance. To learn the semantic and motion association of body parts, we introduce Inter-Semantic Kinematic Topology in Section 3.1, aligning human semantic and kinematic topology with 3D humans to capture complex geometric features of body parts and kinematic correlations. To learn relationships in appearance within body parts, we introduce Intro-Surface Disentanglement in Section 3.2, which disentangles high-frequency features from human body features in each body part, refining the local structures of significant discrepancies on the human surface.
Figure 2: HUGS framework. We introduce Human Gaussian Control with Hierarchical Semantic Graphs (HUGS) as a method for generating Gaussian humans, ensuring both realistic human appearance and anatomical structure. The input initialized point cloud is mapped using the SKT module to establish a graph structure that encodes the semantic topological relationships between different body parts. High-frequency regions are identified by SD for each body part, and the density of Gaussian points is increased in these areas. Finally, the adjusted Gaussian points are rendered to produce the image output.
@@ 3.1 Preliminary
SMPL [28]: The SMPL model is a pre-trained parametric human model representing body shape and pose. In SMPL, body shape and pose are controlled by two main parameters: $\beta$ and $\theta$, where $\beta$ represents 10 parameters for proportions such as height, weight, and head-to-body ratio, and $\theta$ represents 75 parameters for overall body motion posture and relative angles of 24 joints. These parameters enable the SMPL model to generate various human body shapes and poses.
@@
3.2 Inter-Semantic Kinematic Topology
In this section, we introduce our proposed Inter-Semantic Kinematic Topology Aware method, which is based on a topological graph of Gaussian point semantic relationships. This method comprises two submodules: Semantic Alignment Gaussian and Inter-Joint Graph. By learning the motion relationships at the connections of different body parts, it addresses the blurring issues at these junctions in 3D reconstructions.
Applying 3D Gaussian Splatting directly to human reconstruction presents several challenges. Relying on 2D pixel-level supervision to refine the initial coarse Gaussian points distribution, this approach lacks explicit learning of Gaussian points based on different body parts of the human anatomy and overlooks the topological relationships between these parts. The human complex motion may bring about local structures being occluded, generating semantic confusion in Gaussian points at the junction of body parts. Gaussian points fail to correctly match their corresponding spatial positions in the human body, resulting in geometric distortion in the occluded parts. Therefore, considering the intrinsic geometric structures of different body parts and their topological relationships is crucial in addressing the reconstruction problem of local geometric distortion in the occluded parts.
@@ 3.2 Inter-Semantic Kinematic Topology
Semantic Alignment Gaussian: The semantic attributes of Gaussian points explicitly inject semantics into the 3D Gaussian optimization process. Utilizing these semantic labels, we establish connections between 2D and 3D spaces. In the 3D space, we expect points that are close in distance to have similar semantic attributes. When rendered onto a 2D plane, these semantic attributes are also projected. We use a pre-trained CLIP model to establish correspondences between semantically segmented 2D supervisory images and the rendered semantic projection masks. In this way, we can better associate Gaussian points belonging to the same part of the body, enrich the 2D supervisory information, and strengthen the connection between the 3D space and 2D supervision. The loss function formula corresponding to semantic consistency constraint is shown as follows:
$\mathcal{L}_{\text{semantic}}=\frac{1}{N}\sum_{i=1}^{N}(D(\mathcal{X}_{i},y_{i})+\frac{1}{k}\sum_{j\in Q_{i}}D(\mathcal{X}_{j},y_{i})),$ (3)
where $N$ represents the total number of Gaussian points, $y_{i}$ represents the semantic attribute of Gaussian point $i$, $Q_{i}$ contains the $k$ nearest neighbors of Gaussian point $i$ in 3D space. $\mathcal{X}_{i}$ represents the semantic information of the pixel rendered by Gaussian point $i$, 3DGS employs point-based rendering techniques ($\alpha$-blending) to render the color attributes of 3D Gaussian points onto the 2D-pixel plane, function $D$($\cdot$) is used to determine whether the Gaussian point semantic properties are consistent with the true semantics corresponding to the pixels to which they are mapped. Similarly, we blend the semantic attributes of the human body from 3D space to a 2D-pixel plane by mixing them in the order of the influence of on pixels.The rendering formula is as follows:
@@
where $\mathcal{X}_{i}$ represents the semantic labels of pixel $i$, which are derived from Gaussian point semantic attribute $\alpha$-blending(Equation 2), $\mathcal{N}$ represents the set of all Gaussian points that overlap with the pixels rendered by Gaussian point $i$, $y_{p}$ represents the semantic attribute corresponding to the 3D Gaussian point $p$, and $\alpha_{p}$ represents the influence factor of the Gaussian point $p$ for rendering pixels.
Inter-Joint Graph Although the semantic consistency constraint can independently optimize the geometric structure of each body part, the different parts of the human body are not completely independent. There are mutual influences and collaborative effects between various body parts. For example, the position and posture of the arms are influenced by the positions and postures of the palms and shoulders. However, due to the complexity of human motion, the linkage effects between body parts cannot be explicitly expressed. Therefore, we explore using the explicit topological structure of the human body to learn the implicit linkage effects between Gaussian points of different body parts. To achieve this goal, we introduce topological coherence constraints to learn the correlations between different body parts. To incorporate topological information into the human reconstruction process, we use the spatial position information of Gaussian points to create a topological graph for the human body, where each Gaussian point corresponds to a node in the graph. For each node, we calculate the $k$ (same in Equation 3) nearest neighbors based on spatial distance and add edges, with the edge weights represented by the distances. In the topological graph, we use random walk algorithms [7, 37, 38, 32] to propagate information and sample, generating positional embedding vector for each node to capture the topological associations between different Gaussian points. Additionally, we establish a standard prior topological relationship graph for 15 key regions of the human body based on segmentation and joint associations as supervisory information, where each key part is associated with the two adjacent parts in the human body topology diagram. We then sample from all point clouds to construct contrastive learning samples. If the semantic labels of the two sampled points match the prior associations, they are considered positive samples; otherwise, they are negative samples. The loss function corresponding to the topological coherence constraint is as follows:
where $N$ represents the total number of Gaussian points, $M$ represents the number of positive and negative samples used for contrast learning, $t_{i}$ represents the embedding vector for node $i$, which contains topological information obtained by the random walk algorithm, $t_{j}^{+}$ represents the positive samples from body parts that have a linkage effect with the body part corresponding to node $i$, such as Arm and Hand, $t_{j}^{-}$ is the negative sample from body parts that do not have a linkage effect with the body part corresponding to node $i$.
@@
Based on Semantic Kinematic Topology, we initially express both the geometric structure of the body parts and the topological correlations between them. However, we observed fuzzy geometric shapes in the local structures with significant discrepancies on the human surface, such as clothing wrinkles and muscle textures. In the next subsection, we will introduce the Intro-Surface Disentanglement module designed to address this issue.
3.3 Intro-Surface Disentanglement
In this section, we present our proposed Intro-Surface Disentanglement method, which is based on the high-frequency extraction between Gaussian points with the same semantics, structured as a graph. This method aims to address the internal blurriness within body parts in 3D reconstructions by learning the appearance relationships within these parts.
Based on Semantic Kinematic Topology, we initially express both the geometric structure of the body parts and the topological correlations between them. However, we observe fuzzy geometric shapes in the local structures with significant discrepancies on the human surface, such as clothing wrinkles and muscle textures. These discrepancies are primarily found in high-frequency regions of the human surface. Therefore, capturing high-frequency features becomes essential to guide the optimization of local structures with significant differences. As a result, we explore disentangling high-frequency features from the global human features and utilize them to refine Gaussian point density and distribution corresponding to the local structures with significant discrepancies.
To address this challenge, we introduce a surface decoupling module. High-frequency signals manifest as anomalous nodes in the human topological graph, which are significantly different from adjacent nodes. First, We cluster all the Gaussian points based on their semantic label vectors. In each iteration, we identify specific nodes as candidate high-frequency nodes by evaluating the average magnitude of structural differences between the selected node and all nodes in the same cluster. Ultimately, we select the nodes with the highest average magnitude of structural differences within each cluster as high-frequency nodes. The algorithm formula is as follows:
$H_{m}=\mathop{\mathrm{argmax}}\limits_{i\in C_{m}}\frac{1}{C_{m}-1}\sum_{j\in C_{m},j\neq i}\text{similar}\left(A_{i},A_{j}\right).$ (6)
where $A_{i}$ represents the basic attribute of Gaussian points (color, opacity, and radii), $C_{m}$ represents the set of all points with label m. $\text{similar}(\cdot)$ is used to evaluate the similarity between two nodes.